OpenStudy (anonymous):

Find the second derivative of the function f(v)=3v^2+4v-5/v

OpenStudy (anonymous):

do you know how to find the first derivative?

OpenStudy (anonymous):

kind of not really

OpenStudy (anonymous):

do i use the quotient rule?

OpenStudy (anonymous):

f'(v)= 6v+4-5v^-2

OpenStudy (anonymous):

then take the derivative of that

OpenStudy (anonymous):

f''(v)= 6-10v^-3

OpenStudy (anonymous):

that's the final answer?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

My teacher gave us the answer an he put -10/v^2

OpenStudy (anonymous):

to the third power, sorry!

OpenStudy (anonymous):

are you sure? because 3v^2 first derivative= 6v second derivative= 6 so it should be 6- that

OpenStudy (anonymous):

I meant = -10/v^3

OpenStudy (anonymous):

6-10/v^3 should be the answer

OpenStudy (anonymous):

Alright thanks!

OpenStudy (anonymous):

no problem!